You are considering investing in a startup company A. Company A is not expected to generate profits in the near future. After careful analysis, you and your colleagues determine that Company A will start generating profits of $400,000 at the end of the 6th year. Thereafter, it will grow at 11% per year forever. Using a discount rate of 18%, and assuming you will be acquiring 50% of the company, what is the amount you would be willing to invest today?
Answer #1

Answer:

Annual Profit A=$4,00,000 at end of 6th year

Discount rate r=18%

Growth rate g=11%

PV of Total profit at end of fifth year P=A/(r-g)=400000/(18%-11%)=$5714285.71

PV of Profit at the beginning of startup=P/(1+r)^5=5714285.71/(1+18%)^5=$2497766.94

Since We are planning to acquire only 50% of the company we should invest 50%*2497766.94=$1248883.47